Circular No. 7524

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2000 SO_{253} (LINEAR) An apparently asteroidal object discovered by LINEAR (MPS 20197,

21370; discovery observation below) has been found to be cometary (highly condensed 5′′ coma and a 15′′ tail in p.a. 45^{o}) on 300-s R-band CCD exposures taken on Nov. 24.3 UT by C. W. Hergenrother and A. E. Gleason with the Steward Observatory 1.54-m reflector. Additional observations and orbital elements (T = 2001 May 2.1 TT, q = 1.694 AU, i = 3^{o}.7, P = 7.04 yr) are given on MPEC 2000-W39.

2000 UT  α_{2000}  δ_{2000}  m_{2}
Sept.24.34851  0^{h}51^{m}19.^{s}53  −0^{o}55′59′′.0  19.8

SUPERNOVAE 2000em, 2000en, 2000eo, 2000ep, 2000eq L.-G. Strolger, University of Michigan and Cerro Tololo Interamerican

Observatory (CTIO), reports the discovery by the Nearby Galaxies Supernova Search team (cf. IAUC 7297, 7416; plus J. C. Seguel, J. Krick, and A. Block) of several supernovae with the Kitt Peak 0.9-m telescope (+ Mosaic camera; limiting mag R ∼ 21). Search frames were compared to template frames obtained a year or more previously.

SN 2000 UT α_{2000} δ_{2000} R Offset
2000em Nov. 16 0^{h}35^{m}29.^{s}82 − 2^{o}39′27′′.4 18.3 4′′.5 W, 2′′.5 N
2000en Nov. 16 0 43 08.37 − 0 23 17.6 20.2 1′′.8 W, 2′′.5 N
2000eo Nov. 16 3 09 08.17 −10 17 55.3 19.2 12′′.0 W, 20′′.9 S
2000ep Nov. 16 9 19 49.86 − 4 43 28.2 18.7 2′′.8 E, 4′′.5 S
2000eq Nov. 17 21 03 57.97 − 9 41 31.8 20.5 7′′.3 W, 7′′.3 N

Direct confirmation of each object was made by P. Candia using the CTIO 0.9-m telescope on Nov. 18 and 19. Spectra (range 400–900 nm) were obtained by R. C. Smith and Strolger using the Mayall 4-m telescope on Nov. 20–22. From the spectra, N. B. Suntzeff and M. M. Phillips identify SN 2000em as most likely a peculiar type-II event at z = 0.019, resembling SN 1997D. The spectra of SN 2000en indicate that it is a type-Ia event, a few days past maximum light. SN 2000eo in MCG –2-9-3 is a type-IIn supernova with Wolf-Rayet-like He II and N III/C III emission, similar to SN 1998S, at z = 0.010. SN 2000ep has spectra consistent with a normal type-II supernova within a few weeks of maximum, at z = 0.024. SN 2000eq appears to be a type-Ia supernova, 2–3 weeks past maximum light.

2000 November 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green